OpenStudy (anonymous):

Suppose you roll a number cube. What is the theoretical probability of rolling a number less than 6? A. 1/3 B. 1/6 C. 2/3 D. 5/6

OpenStudy (anonymous):

d

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

are you sure?

OpenStudy (anonymous):

A dice has 6 sides, 1-6 and 1-5 is less than 6 also the rule of probability states p(x)+p(y)=1 therefore if p(x)=6/6 therefore p(y)= 1-6/6 p(y)=5/6

OpenStudy (anonymous):

and yeah i'm pretty sure :)